puppylover1234:

Eloy surveyed 700 students at his middle school and asked them to name their favorite color. The results of the survey showed that 63 students named yellow as their favorite color. What percentage of students in the survey named yellow as their favorite color? A)63% B)0.09% C)11.11% D)9%

puppylover1234:

@MrMudd183

MrMudd183:

What do you think the first thing to do is?

puppylover1234:

Multiply?

puppylover1234:

or divide?

MrMudd183:

The first thing I would do is divide 700 by 100% and 63 by 100%

MrMudd183:

That will turn both into percents

MrMudd183:

Not quiet, 700 divided by 100% is 7 Basically when you divide by 100, you're moving the decimal place to the left twice

puppylover1234:

oh well im dumb

puppylover1234:

thank you tho!

MrMudd183:

You're welocme. So what do you think the correct answer choice would be after you divide?

puppylover1234:

63%?

MrMudd183:

Correct, good job! C:

puppylover1234:

Thank you!!!!!

MrMudd183:

For example, if you do 12 divided by 100%...it'll give you 0.12, which would become 12%

supie:

I believe your answer is incorrect mudd, a precent is out of 100, not 700

supie:

right...?

supie:

you would have \(\frac{63}{700}\) of those who chose yellow so we need to make 63/700 as a precent the way you would do that is to get the denominator to be 100 which means you would have too do \(\frac{63}{700} \div \frac{100}{100}=\frac?? \) divide straight across

supie:

Yes...?

MrMudd183:

Yeah, well I was thinking of a different method that I do in my head and works for me

supie:

yeah I think I messed up too dont do 'which means you would have too do \(\frac{63}{700} \div \frac{100}{100}=\frac?? \) divide straight across'

supie:

but yeah a precent it out of 100 so you would have to get 63/700 as a precent

MrMudd183:

If you combine what me and supie are saying, I'm thinking it would be around 0.63/7.00 or something like that

supie:

uh no to find 63/700 as a precent we would have to do \(63\div700\times100\)

puppylover1234:

oh okay, I thought it would be like .63 but that's not an answer choice :)

supie:

do you understand @MrMudd183?

MrMudd183:

Yeah, i'm just trying to get some work for me done

supie:

Alright, what did you get when you did 63÷700×100 ? @puppylover1234

puppylover1234:

9

supie:

Yes, correct! your answer is 9,

puppylover1234:

oh okay thank you!!!!

supie:

np :)

puppylover1234:

if i have any more problems can i tag you?

puppylover1234:

;-; I've been broke so many times I don't know what to believe ;-;

Laylalyssa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @puppylover1234 if i have any more problems can i tag you? \(\color{#0cbb34}{\text{End of Quote}}\) feel free to tag supie

puppylover1234:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Laylalyssa \(\color{#0cbb34}{\text{Originally Posted by}}\) @puppylover1234 if i have any more problems can i tag you? \(\color{#0cbb34}{\text{End of Quote}}\) feel free to tag supie \(\color{#0cbb34}{\text{End of Quote}}\) thank you so much ;-;

Laylalyssa:

anytime~!